What is the minimum energy of a thermal oscillator in a blackbody producing radiation at λmax at a temperature of 212 °F?
Answer #1

Converting temperature to Kelvin

T=(212+459.67)*(5/9)=373.15 K

From Wien's displacement law

the minimum energy of a thermal oscillator in a blackbody

E=2.56*10-20J or 0.16 eV